# Swin fMRI Transformer Predicts Early Neurodevelopmental Outcomes from Neonatal fMRI

### v1

@@ Abstract
Brain development in the first few months of human life is a critical phase characterized by rapid structural growth and functional organization. Accurately predicting developmental outcomes during this time is crucial for identifying delays and enabling timely interventions. This study introduces the SwiFT (Swin 4D fMRI Transformer) model, designed to predict Bayley-III composite scores using neonatal fMRI data from the Developing Human Connectome Project (dHCP). To enhance predictive accuracy, we apply dimensionality reduction via group independent component analysis (ICA) and pretrain SwiFT on large adult fMRI datasets to address the challenges of limited neonatal data. Our analysis shows that SwiFT significantly outperforms baseline models in predicting cognitive, motor, and language outcomes, leveraging both single-label and multi-label prediction strategies. The model’s attention-based architecture processes spatiotemporal data end-to-end, delivering superior predictive performance. Additionally, we use Integrated Gradients with Smoothgrad sQuare (IG-SQ) to interpret predictions, identifying neural spatial representations linked to early cognitive and behavioral development. These findings underscore the potential of Transformer models to advance neurodevelopmental research and clinical practice.
@@ 1 Introduction
Brain development during the first few months of life is a period of rapid structural and functional reorganization, making it a critical window for identifying potential neurodevelopmental deficits. Accurately predicting developmental outcomes during this period is essential for enabling early interventions that can mitigate the lifelong impact of developmental delays. Neonatal fMRI data, such as those from the Developing Human Connectome Project (dHCP), have shown potential for predicting neurodevelopmental outcomes (Li et al. 2024) (Xia et al. 2024) (Gondová et al. 2023) (Eyre et al. 2021). However, the spatiotemporal complexity of neonatal brain activity presents significant challenges for conventional analysis methods.
This study investigates the potential of the Swin 4D fMRI Transformer (SwiFT) (Kim et al. 2023), a deep learning architecture designed to process high-dimensional fMRI data, for predicting neurodevelopmental outcomes from neonatal fMRI. Specifically, we aim to predict Bayley Scales of Infant and Toddler Development, Third Edition (Bayley-III/BSID-III) composite scores, encompassing cognitive, language, and motor skills, using neonatal fMRI from the dHCP dataset. To address the challenges of limited neonatal data and high dimensionality, we explore dimensionality reduction using group Independent Component Analysis (ICA) and pretraining SwiFT on large publicly available adult fMRI datasets.
We hypothesize that (1) SwiFT will outperform existing baseline models in predicting Bayley-III scores due to its ability to capture complex spatiotemporal patterns in fMRI data and (2) incorporating ICA features and pretraining will further improve predictive accuracy. We further aim to interpret SwiFT’s predictions using Integrated Gradients with Smoothgrad sQuare (IG-SQ) (Rahman et al. 2022) (Adebayo et al. 2018) (Kokhlikyan et al. 2020) to identify brain regions associated with early cognitive, language, and motor development.
This research advances the state of neurodevelopmental prediction by introducing a robust and interpretable framework for neonatal fMRI analysis. By addressing key methodological challenges, our findings lay the groundwork for future applications in neurodevelopmental research and clinical practice.
@@
In Section 2, we introduce the dataset, its acquisition, (additional) processing and target values. We also explain the SwiFT Transformer, our motivations and how we will interpret the results of the model. In Section 3, we will explain in what settings we perform our experiments, and introduce (i) the baselines, (ii) SwiFT running on (pre-processed) raw fMRI data with single-label and multi-label prediction, and (iii) SwiFT running on group-ICA-extracted features with single-label and multi-label prediction. In the end, we give an overview of the performance gain through our methods. In Section 4, we perform IG-SQ on the best performing model and provide our interpretation of the IG maps on risk of delay in cognitive, language and motor skills. Moreover, we provide a small glimpse into what model interpretations on brain-network-level may look like by using ICA-extracted features. Lastly, we conclude our research in Section 5 through discussing future directions and limitations of our current work.
@@ 2 Methods
This study focuses on Bayley Scales of Infant and Toddler Development, Third Edition (Bayley-III), composite scores in cognitive, language, and motor domains. Developmental delays are categorized based on established thresholds  (Johnson, Moore, and Marlow 2014):
@@
Following prior studies  (He et al. 2018) (He et al. 2020a), we simplify the classification task into a binary format, with a threshold of 85 distinguishing low- and high-risk cases. Regression tasks are also conducted to predict continuous Bayley-III scores, providing finer-grained insights into developmental outcomes. By integrating these approaches, we aim to enhance the accuracy and interpretability of predictions using SwiFT.
@@
The Developing Human Connectome Project (dHCP) dataset provides a rich and representative basis for developing predictive models of early neurodevelopment. It includes 783 neonates born between 23–44 gestational weeks, with imaging performed between 26–45 post-menstrual weeks. Resting-state fMRI (rs-fMRI) data were acquired using a 3T Philips Achieva scanner and preprocessed via the dHCP pipeline for motion correction, denoising, and spatial normalization.
For this study, Bayley-III scores in cognition, motor, and language domains were available for 739 infants, with 619 of them also having usable fMRI data. The composite scores showed a mean close to 100, with imbalances in developmental delay cases: $\approx 5\%$ for motor and cognitive domains and $\approx 18\%$ for language. Strong positive correlations were observed between cognitive, motor, and language scores (63.08%, 57.27%, and 51.94%, respectively). Maternal factors (e.g., smoking) showed no significant correlation with scores.
Figure 1: Distributions of composite scores based on binary classification. Assumptions made in (Johnson, Moore, and Marlow 2014) largely hold true.
@@
MRI data were acquired on a 3T Philips Achieva system with a neonatal 32-channel head coil. The imaging protocol was tailored to the unique characteristics of the neonatal brain, which has different tissue properties compared to the adult brain, including higher water content and longer T1 and T2 relaxation times. Resting-state fMRI (rs-fMRI) was captured with a multiband-accelerated echo-planar imaging sequence (TR: 392 ms, TE: 38 ms, resolution: 2.15 mm³) over roughly 15 minutes (2300 volumes). Processing followed the dHCP pipeline, which is specifically designed to address the challenges of neonatal imaging. This included motion and distortion correction, skull stripping, ICA-based denoising with FIX, and high-pass temporal filtering to remove slow drifts.
To align the data to a neonatal-specific template space, we adopted a pipeline inspired by the automated resting-state functional framework developed for the dHCP dataset (Fitzgibbon et al. 2020). Specifically, we utilized the 40-week T1-weighted template from the extended dHCP Augmented Volumetric Atlas (Schuh et al. 2018), downsampled from a spatial resolution of 0.5 $mm^{3}$ isotropic to 2.15 $mm^{3}$ isotropic to match the native resolution of the fMRI images. Image registration was performed using ANTs (Tustison et al. 2021) on the mean image of all timepoints, with the resulting transformation applied independently to each timeframe in parallel (Tange 2023). The processed outputs served as inputs for the subsequent experiments.
@@
To predict neurodevelopmental outcomes from neonatal fMRI, we employed SwiFT, a Transformer-based architecture designed to model high-dimensional spatiotemporal fMRI data. SwiFT processes raw 4D fMRI volumes, capturing local and global dependencies critical for neurodevelopmental predictions. The model operates by dividing 4D fMRI data into patches, embedding them as tokens, and processing these tokens through multiple layers of 4D Swin Transformer blocks. These blocks utilize a shifted window multi-head self-attention mechanism, enabling efficient modeling of complex spatiotemporal patterns while reducing computational complexity through patch merging. Previously, SwiFT demonstrated robust performance on diverse fMRI tasks, including sex classification (97.7% accuracy on HCP) and age prediction (MAE of 3.40 years on UK Biobank), showcasing its effectiveness in capturing brain dynamics across populations.
@@ 3.1 Experimental Settings
All experiments were conducted on the Perlmutter supercomputer, an HPE Cray EX system. A holdout strategy with a 70:15:15 data split was used for training, validation, and testing. To address dataset imbalance (see Section 2), stratified splits were applied based on BSID-III scores. For benchmarking, stratified 5-fold cross-validation ensured consistent and fair splits across experiments.
For binary classification, we used weighted focal loss (Lin et al. 2018) to prioritize harder-to-classify (at-risk) cases. This approach significantly improved performance, especially for these challenging instances.
Following the original SwiFT paper (Kim et al. 2023), we padded input images to 64x64x64 to align with the model architecture, adapting from the 96x96x96 padding used in the original work due to our native image size of 45x55x45.
For classification tasks, we evaluated performance using the area under the ROC curve (AUC), accuracy, and balanced accuracy (ACCbal) to account for class imbalance. For regression tasks, mean absolute error (MAE) and mean squared error (MSE) were used, along with adjusted metrics (MAEadj and MSEadj) to reflect re-scaled values for interpretability.
@@ 3.2 Baselines
We established baseline performance using ROI-based deep learning models, including BrainNetCNN (Kawahara et al. 2017), VanillaTF, and Brain Network Transformer (BNT) (Kan et al. 2022), which take functional connectivity matrices as input (computed via Pearson correlations of 87 brain regions in the dHCP neonatal brain atlas). Hyper-parameters and implementations followed (Kan et al. 2022). Additionally, we included XGBoost (Chen and Guestrin 2016) as a traditional machine learning baseline, using the flattened upper triangular correlation matrix as input.
Table 1: Mean regression performance of each baseline model with standard deviation through 5-fold cross-validation
@@
Table 2: Mean classification performance of each baseline model with standard deviation through 5-fold cross-validation.
@@
We evaluated the SwiFT model’s performance on fMRI data in an end-to-end manner, focusing on single-label and multi-label prediction tasks. Multi-label prediction was hypothesized to enhance performance due to strong correlations among Bayley scores (see Section 2), as supported by prior work (He et al. 2020b).
The models were fine-tuned across a predefined hyperparameter search space, including data augmentation strategies (affine-only, intensity-only, and combined transformations), learning rates, batch sizes, and patch sizes. Sequence length was varied to evaluate its impact: a baseline length of 20 was used per the original SwiFT paper (Kim et al. 2023), with additional experiments using lengths of 50 and 100 to assess temporal dynamics. Longer sequences have shown promise in prior studies for predicting intelligence in young adults and elders (Kim et al. 2023).
To leverage publicly available fMRI datasets, we pre-trained SwiFT using contrastive self-supervised learning (Dave et al. 2022) on data from the UK Biobank (UKB)(Alfaro-Almagro et al. 2018b; Miller et al. 2016a), Human Connectome Project (HCP)(Van Essen et al. 2013), and Adolescent Brain Cognitive Development (ABCD) study(Casey et al. 2018). While this approach utilizes larger datasets, differences between neonatal and adult or adolescent brain data present potential limitations. To address size discrepancies, neonatal images were padded to 96x96x96, contrasting with experiments trained from scratch, where images were padded to 64x64x64 (see Section 3.1).
Figure 2: Rationale of using large amounts of brain fMRI data across different datasets to make up for the little amount of data available for neonates.
@@
The choice of sequence length had a notable impact on the model’s performance. A length of 50 yielded the best results, while shorter (20) or longer (100) sequences performed slightly worse, suggesting that an optimal balance is necessary.
As shown in Table 3, pre-training resulted in only minor improvements in predictive power for BSID-III scores. However, increased padding might have negatively influenced performance. Compared to baseline models, SwiFT provided significant performance gains for both regression and classification tasks. For cognitive and language predictions, adjusted MAE improved by approximately 11%, while gains for motor regression were minimal. In classification, balanced accuracy ($\text{ACC}_{bal}$) improved by 5-7% points, exceeding guessing levels.
Table 3: Mean performance with standard deviation through 5-cross-validation by using raw fMRI, facilitating single-label predictions.
@@
Multi-Label Prediction
To predict all three classes simultaneously, we used a shared classification/regression head, a design choice based on the expected correlation among these values (see Section 2).
Sequence length changes showed similar effects to single-label predictions. As shown in Table 4, pre-training provided only marginal improvements, highlighting the limited transferability of adult data to neonatal predictions. However, multi-label regression results align with previous studies (He et al. 2020b), showing performance gains across all classes for both regression and classification. These improvements may stem from shared learning across developmental domains, enabling the model to capture complex, interrelated features of early brain development.
Table 4: Mean performance with standard deviation through 5-cross-validation by using raw fMRI, facilitating multi-label predictions.
@@ 3.4 Dimensionality Reduction via Group ICA
Group Independent Component Analysis (Group ICA) (Hyvärinen 1999; Beckmann and Smith 2004) is a widely used data-driven technique in neuroimaging that decomposes fMRI data into independent spatial and temporal components. These components represent distinct neural processes or functional networks. Group ICA reduces the dimensionality of high-dimensional fMRI data while preserving significant information and isolating biologically interpretable brain networks.
In this study, we use Group ICA features as model inputs for both single-label and multi-label learning. These features encapsulate complex brain activity patterns in a compact representation, enhancing expressiveness and facilitating brain network-level analysis. We hypothesize that this approach will improve model interpretability and predictive performance by leveraging the reduced dimensionality and biologically relevant features.
Preparation
@@
To generate the Group ICA map, we selected a subset of normally developing neonates, defined as those with cognitive, language, and motor BSID-III scores above 85, indicating low risk of developmental delay. Only normally developing data were used to reduce variability and noise, ensuring a more reliable feature set.
Out of 725 assessed subjects, 431 met the criteria for normal development, with 358 having usable fMRI data. To further enhance stability, we excluded neonates with gestational ages outside 36 to 44 weeks, as defined by the dHCP volumetric atlas (Schuh et al. 2018), resulting in a final cohort of 348 subjects. From this group, we randomly selected 100 neonates for Group ICA analysis.
@@
The workflow for Group ICA and functional connectivity analysis (see Figure 3) follows methodologies established for the UKB dataset (Alfaro-Almagro et al. 2018a; Miller et al. 2016b). For the selected 100 subjects, we used MIGP (MELODIC’s Incremental Group-PCA) to extract the top 1200 weighted spatial eigenvectors, approximating traditional PCA while remaining computationally efficient for large datasets (Smith et al. 2014). Spatial Independent Component Analysis (spatial-ICA) was then applied using FSL MELODIC (Hyvärinen 1999; Beckmann and Smith 2004) to parcellate the data into spatially independent components (ICs). We evaluated dimensionalities of 25 and 100 ICs, consistent with prior work on the UKB dataset (Alfaro-Almagro et al. 2018a; Miller et al. 2016b; Smith et al. 2013, 2009). Higher dimensionalities result in smaller, more numerous regions within spatial maps. Since FIX denoising was performed during preprocessing, we did not identify artifactual components, assuming their removal.
Using the generated group ICA maps, we performed dual regression to obtain subject-specific spatial maps for each IC. Dual regression comprises two steps: (i) spatial regression, projecting group-level IC maps onto individual data, and (ii) temporal regression, deriving subject-specific timeseries for each IC. To further explore functional connectivity, we conducted seed-to-voxel (seed2voxel) analysis, correlating timeseries of specific brain regions with those of all other voxels. The resulting whole-brain functional connectivity maps served as input features for the SwiFT model to predict neurodevelopmental outcomes.
Figure 3: UML Flowchart depicting the workflow to extract features and connectivity maps by leveraging Group Independent Component Analysis, heavily inspired by (Alfaro-Almagro et al. 2018a)(Miller et al. 2016b) and (Gal et al. 2022).
Results

### v2

# %****␣bookchapter.tex␣Line␣75␣****https://informatics.tuwien.ac.at/

@@ Abstract
Brain development in the first few months of human life is a critical phase characterized by rapid structural growth and functional organization. Accurately predicting developmental outcomes during this time is crucial for identifying delays and enabling timely interventions. This study introduces the SwiFT (Swin 4D fMRI Transformer) model, designed to predict Bayley-III composite scores using neonatal fMRI from the Developing Human Connectome Project (dHCP). To enhance predictive accuracy, we apply dimensionality reduction via group independent component analysis (ICA) and pretrain SwiFT on large adult fMRI datasets to address the challenges of limited neonatal data. Our analysis shows that SwiFT significantly outperforms baseline models in predicting cognitive, motor, and language outcomes, leveraging both single-label and multi-label prediction strategies. The model’s attention-based architecture processes spatiotemporal data end-to-end, delivering superior predictive performance. Additionally, we use Integrated Gradients with Smoothgrad sQuare (IG-SQ) to interpret predictions, identifying neural spatial representations linked to early cognitive and behavioral development. These findings underscore the potential of Transformer models to advance neurodevelopmental research and clinical practice.
Keywords:
fMRI Transformer, Developing Human Connectome Project, Bayley Scales of Infant Development, Personalized Therapy, XAI
@@ 1 Introduction
Brain development during the first few months of life is a period of rapid structural and functional reorganization, making it a critical window for identifying potential neurodevelopmental deficits. Accurately predicting developmental outcomes during this period is essential for enabling early interventions that can mitigate the lifelong impact of developmental delays. Neonatal fMRI data, such as those from the Developing Human Connectome Project (dHCP), have shown potential for predicting neurodevelopmental outcomes [1] [2] [3] [4]. However, the spatiotemporal complexity of neonatal brain activity presents significant challenges for conventional analysis methods. This study investigates the potential of the Swin 4D fMRI Transformer (SwiFT) [19], a deep learning architecture designed to process high-dimensional fMRI data, for predicting neurodevelopmental outcomes from neonatal fMRI. Unlike existing methods, SwiFT leverages 4D spatiotemporal attention mechanisms to effectively capture dynamic brain connectivity patterns, offering a novel approach for analyzing neonatal fMRI data. Specifically, we aim to predict Bayley Scales of Infant and Toddler Development, Third Edition (Bayley-III/BSID-III) composite scores, encompassing cognitive, language, and motor skills, using neonatal fMRI from the dHCP dataset. To address the challenges of limited neonatal data and high dimensionality, we explore dimensionality reduction using group Independent Component Analysis (ICA) and pretraining SwiFT on large publicly available adult fMRI datasets. We hypothesize that (1) SwiFT outperforms baseline models in predicting Bayley-III scores due to its ability to capture complex spatiotemporal patterns in fMRI data and (2) incorporating ICA features and pretraining further improves predictive accuracy. Furthermore, we aim to interpret SwiFT’s predictions using Integrated Gradients with Smoothgrad sQuare (IG-SQ) [28] [29] [30] to identify brain regions associated with early cognitive, language, and motor development. This research advances the state of neurodevelopmental prediction by introducing a robust and interpretable framework for neonatal fMRI analysis. By addressing key methodological challenges, our findings lay the groundwork for future applications in neurodevelopmental research and clinical practice.
@@
In Section 2, we introduce the dataset, its acquisition, (additional) processing and target values. We also explain SwiFT, our motivations and how we will interpret the results of the model. In Section 3, we will explain in what settings we perform our experiments, and introduce (i) the baselines, (ii) SwiFT running on (pre-processed) raw fMRI data with single-label and multi-label prediction, and (iii) SwiFT running on group-ICA-extracted features with single-label and multi-label prediction. In the end, we give an overview of the performance gain through our methods. In Section 4, we perform IG-SQ on the best performing model and provide our interpretation of the IG maps on risk of delay in cognitive, language and motor skills. Moreover, we provide a small glimpse into what model interpretations on brain-network-level may look like by using ICA-extracted features. Lastly, we conclude our research in Section 5 through discussing future directions and limitations of our current work.
@@ 2 Methods
This study focuses on Bayley Scales of Infant and Toddler Development, Third Edition (Bayley-III), composite scores in cognitive, language, and motor domains. Developmental delays are categorized based on established thresholds  [8]:
@@
Following prior studies  [10] [11], we simplify the classification task into a binary format, with a threshold of 85 distinguishing low- and high-risk cases. Regression tasks are also conducted to predict continuous Bayley-III scores.
@@
[…]
Figure 1: Distributions of composite scores based on binary classification. Assumptions made in [8] largely hold true.
@@
MRI data were acquired on a 3T Philips Achieva system with a neonatal 32-channel head coil. The imaging protocol was tailored to the unique characteristics of the neonatal brain. Resting-state fMRI (rs-fMRI) was captured with a multiband-accelerated echo-planar imaging sequence (TR: 392 ms, TE: 38 ms, resolution: $2.15\text{mm}^{3}$) over 15 minutes (2300 volumes). Processing followed the dHCP pipeline, which is designed to address the challenges of neonatal imaging. This included motion and distortion correction, skull stripping, ICA-based denoising with FIX, and high-pass temporal filtering to remove slow drifts. To align the data to a neonatal-specific template space, we adopted a pipeline inspired by the automated resting-state functional framework developed for the dHCP dataset [5]. Specifically, we utilized the 40-week T1-weighted template from the extended dHCP Augmented Volumetric Atlas [6], downsampled from a spatial resolution of 0.5 $mm^{3}$ isotropic to 2.15 $mm^{3}$ isotropic to match the native resolution of the fMRI images. Image registration was performed using ANTs [7] on the mean image of all timepoints, with the resulting transformation applied independently to each timeframe in parallel. The processed outputs served as inputs for the subsequent experiments.
@@
To predict neurodevelopmental outcomes from neonatal fMRI, we employed SwiFT, a Transformer-based architecture designed to model high-dimensional spatiotemporal fMRI data. SwiFT processes raw 4D fMRI volumes, capturing local and global dependencies critical for neurodevelopmental predictions. The model operates by dividing 4D fMRI data into patches, embedding them as tokens, and processing these tokens through multiple layers of 4D Swin Transformer blocks. These blocks utilize a shifted window multi-head self-attention mechanism, enabling efficient modeling of complex spatiotemporal patterns while reducing computational complexity through patch merging. Previously, SwiFT demonstrated robust performance on diverse fMRI tasks, including sex classification (97.7% accuracy on HCP) and age prediction (MAE of 3.40 years on UK Biobank), showcasing its effectiveness in capturing brain dynamics across populations [19].
@@ 3.1 Experimental Settings
All experiments were carried out on the Perlmutter supercomputer, an HPE Cray EX system. A holdout strategy with a 70:15:15 data split was used for training, validation, and testing. To address data imbalance (see Section 2), stratified splits were applied based on the BSID-III scores. For benchmarking, 5-fold cross-validation ensured consistent and fair splits between experiments. For binary classification, we used weighted focal loss [26] to prioritize harder-to-classify (at-risk) cases. This approach significantly improved performance, especially for these challenging instances. Following the original SwiFT paper [19], we padded input images to 64x64x64 to align with the model architecture, adapting from the 96x96x96 padding used in the original work due to our native image size of 45x55x45. For classification tasks, we evaluated performance using the area under the ROC curve (AUC), accuracy, and balanced accuracy (ACCbal) to account for class imbalance. For regression tasks, mean absolute error (MAE) and mean squared error (MSE) were used, along with adjusted metrics (MAEadj and MSEadj) to reflect re-scaled values for interpretability.
@@ 3.2 Baselines
We established baseline performance using ROI-based deep learning models, including BrainNetCNN [9], VanillaTF, and Brain Network Transformer (BNT) [38], which take functional connectivity matrices as input (computed via Pearson correlations of 87 brain regions in the dHCP neonatal brain atlas). Hyper-parameters and implementations followed [38]. Additionally, we included XGBoost [39] as a traditional machine learning baseline, using the flattened upper triangular correlation matrix as input. Experiments for both classification and regression can be seen in Table 1.
Table 1: Baseline performances on regression (left) and classification (right).
@@
We evaluated the SwiFT model’s performance on fMRI data in an end-to-end manner, focusing on single-label and multi-label prediction tasks. Multi-label prediction was hypothesized to enhance performance due to strong correlations among Bayley scores (see Section 2), as supported by prior work [27]. The models were fine-tuned across a predefined hyperparameter search space, including data augmentation strategies (affine-only, intensity-only, and combined transformations), learning rates, batch sizes, and patch sizes. Sequence length of the input was varied to evaluate its impact: a baseline length of 20 was used per the original SwiFT paper [19], with additional experiments using lengths of 50 and 100 to assess temporal dynamics. Longer sequences have shown promise in prior studies for predicting intelligence in young adults and elders [19]. To leverage publicly available fMRI datasets, we pre-trained SwiFT using contrastive self-supervised learning [24] on data from the UK Biobank (UKB)[23] [22], Human Connectome Project (HCP)[20], and Adolescent Brain Cognitive Development (ABCD) study[21]. While this approach utilizes larger datasets, differences between neonatal and adult or adolescent brain data present potential limitations. To address size discrepancies, neonatal images were padded to 96x96x96, contrasting with experiments trained from scratch, where images were padded to 64x64x64 (see Section 3.1).
Figure 2: Rationale of using large amounts of brain fMRI data across different datasets to make up for the little amount of data available for neonates. We expect that learned features from elderly, adult and child data will generalize to limited neonatal fMRI data and thus improve downstream performance for neonates.
@@
The choice of sequence length had a notable impact on the model’s performance. A length of 50 yielded the best results, while shorter (20) or longer (100) sequences performed slightly worse, suggesting that an optimal balance is necessary. As shown in Table 2, pre-training resulted in marginal improvements in predictive power for BSID-III scores. However, increased padding might have negatively influenced the performance. Compared to baseline models, SwiFT yielded significant performance gains for both regression and classification tasks. For cognitive and language predictions, adjusted MAE improved by approximately 11%, while gains for motor regression were minimal. In classification, balanced accuracy ($\text{ACC}_{bal}$) improved by 5-7% points, exceeding guessing levels.
Table 2: SwiFT’s performance with raw fMRI as input on single-label (left) and multi-label (right) predictions.
@@
Multi-Label Prediction
To predict all three classes simultaneously, we used a shared classification/regression head, a design choice based on the expected correlation among these values (see Section 2). Changes in sequence length showed similar effects to single-label predictions. As shown in Table 2, pre-training provided marginal improvements, highlighting the limited transferability of adult data to neonatal predictions. However, multi-label regression results align with previous studies [27], showing performance gains across all classes for both regression and classification. These improvements may stem from shared learning across developmental domains, enabling the model to capture complex, interrelated features of early brain development.
@@ 3.4 Dimensionality Reduction via Group ICA
Group Independent Component Analysis (Group ICA) [14] [15] is a widely used data-driven technique in neuroimaging that decomposes fMRI data into independent spatial and temporal components. These components represent distinct neural processes or functional networks. Group ICA reduces the dimensionality of high-dimensional fMRI data while preserving significant information and isolating biologically interpretable brain networks. In this study, we use Group ICA features as model inputs for both single-label and multi-label learning. These features encapsulate complex brain activity patterns in a compact representation, enhancing expressiveness and facilitating brain network-level analysis. We hypothesize that this approach will improve model interpretability and predictive performance by leveraging the reduced dimensionality and biologically relevant features.
3.4.1 Preparation
@@
To generate the Group ICA map, we selected a subset of normally developing neonates, defined as those with cognitive, language, and motor BSID-III scores above 85, indicating low risk of developmental delay. Only normally developing data were used to reduce variability and noise, ensuring a more reliable feature set. Out of 725 assessed subjects, 431 met the criteria for normal development, with 358 having usable fMRI data. To further enhance stability, we excluded neonates with gestational ages outside 36 to 44 weeks, as defined by the dHCP volumetric atlas [6], resulting in a final cohort of 348 subjects. From this group, we randomly selected 100 neonates for Group ICA analysis.
@@
The workflow for Group ICA and functional connectivity analysis (see Figure 3) follows methodologies established for the UKB dataset [13] [12]. For the selected 100 subjects, we used MIGP (MELODIC’s Incremental Group-PCA) to extract the top 1200 weighted spatial eigenvectors, approximating traditional PCA while remaining computationally efficient for large datasets [16]. Spatial Independent Component Analysis (spatial-ICA) was then applied using FSL MELODIC [14] [15] to parcellate the data into spatially independent components (ICs). We evaluated dimensionalities of 25 and 100 ICs, consistent with prior work on the UKB dataset [13] [12] [17] [18]. Higher dimensionalities result in smaller, more numerous regions within spatial maps. Since FIX denoising was performed during preprocessing, we did not identify artifactual components, assuming their removal. Using the generated group ICA maps, we performed dual regression to obtain subject-specific spatial maps for each IC. Dual regression comprises two steps: (i) spatial regression, projecting group-level IC maps onto individual data, and (ii) temporal regression, deriving subject-specific timeseries for each IC. To further explore functional connectivity, we conducted seed-to-voxel (seed2voxel) analysis, correlating timeseries of specific brain regions with those of all other voxels. The resulting whole-brain functional connectivity maps served as input features for the SwiFT model to predict neurodevelopmental outcomes.
Figure 3: UML Flowchart depicting the workflow to extract features and connectivity maps by leveraging Group Independent Component Analysis, heavily inspired by [13][12] and [25].
3.4.2 Results